# BabyLM Turns 4: Call for Papers for the 2026 BabyLM Workshop https://babylm.github.io/

### v1

@@ Abstract
BabyLM aims to dissolve the boundaries between cognitive modeling and language modeling. We call for both workshop papers and for researchers to join the 4th BabyLM competition. As in previous years, we call for participants in the data-efficient pretraining challenge in the general track. This year, we also offer a new track: Multilingual.
We also call for papers outside the competition in any relevant areas. These include training efficiency, cognitively plausible research, weak model evaluation, and more.
@@ 1 Introduction: BabyLM
The goals of BabyLM are to bring together multiple disciplines to answer an enduring question: how can a computational system learn language from limited inputs? Cognitive scientists investigate this question by trying to understand how humans learn their native language during childhood. Computer scientists tackle this question by attempting to build efficient machine-learning systems to accomplish this task. BabyLM aims to bridge these communities, encouraging the integration insights from cognitive science into the design of more sample-efficient language models, while also using advances in language modeling architectures to generate new hypotheses and experimental paradigms for cognitive science.
@@ 3.1 Topics
  • Second language acquisition, bilingualism or multilingualism.
@@
The workshop will have a a theme this year: Going beyond English. Previous iterations of BabyLM have primarily focused on English, and with the introduction of the new Multilingual track we aim to inspire workshop submissions to investigate BabyLM-related topics with a focus on other languages. We hope the BabyBabelLM dataset can serve as a starting point for this, but also encourage submissions to introduce new resources that will foster progress on data-efficient language modeling across diverse linguistic settings. In particular, we welcome work that expands coverage to underrepresented and low-resource languages, explores cross-lingual transfer and multilingual training strategies, and examines how typological diversity, morphology, and script variation shape learnability under BabyLM constraints. We are also interested in benchmarks, evaluation protocols, and ethical data practices that support inclusive and culturally grounded language technologies.
@@ 4.2 Training Requirements
Similar to last year year, we require the submission of results for models trained on a fixed amount of input (counting repeated exposures), in particular after at most 100M words for the Strict-small track and after at most 1B words for all other tracks. In most cases, this will mean after 10 epochs on the standard BabyLM corpora. However, because what counts as an epoch may differ across submissions, we instead quantify training by the number of whitespace-separated input words. For the Multilingual track, this value should be computed while taking each language’s Byte Premium into account (Arnett et al., 2024). While participants are welcome to train for longer and report this in their paper, we will only include the model checkpoints that follow these limitations in the leaderboard. Note that it is also allowed to submit a model that is trained on less data: the 100M and 1B word limits are an upper bound on data exposure. For the Strict and Strict-small tracks, the number of words seen by the submission model is considered to be the sum of the number of input words and generated tokens.
@@ 4.3 Updated Dataset
CHILDES (MacWhinney, 2000) Child-directed speech 15M 29M
British National Corpus (BNC), dialogue portion Dialogue 4M 8M
Project Gutenberg (children’s stories) (Gerlach and Font-Clos, 2018) Written English 13M 26M
OpenSubtitles (Lison and Tiedemann, 2016) Movie subtitles 10M 20M
Simple English Wikipedia Written Simple English 7M 15M
Switchboard Dialog Act Corpus (Stolcke et al., 2000) Dialogue <<1M 1M
@@
Recent work from last year’s Interaction Track found that BabyLMs trained on the provided Strict corpus can produce toxic generations and hate speech, which was a proven challenge for sustainable multi-turn generations between a BabyLM and a Teacher model (Salhan et al., 2025). Subsequent analysis by Trhlik et al. (2026) find that the BabyLM corpus is more toxic and hateful than training corpora used in larger models such as BERT, with a comparatively higher presence of toxic and hateful sentences, even containing racist and sexualised training data, contrary to its child-aligned disposition.
@@ 4.4 Evaluation
As in previous years, we will distribute an open-source evaluation pipeline, building on top of our repository from the 2025 challenge. Much of the evaluation will continue to be based on zero-shot probability comparisons of two text sequences. This year’s new track, Multilingual, will follow a similar logic to the existing evaluation for Strict. Specifically, participants will be evaluating their trained models on a combination of zero-shot and finetuning-based tasks for English, Dutch and Chinese. More details about the evaluation pipeline and the set of tasks will be released subsequently. Similar to previous years, we will release a set of hidden tasks shortly before the submission deadline, to test the robustness of models on previously unseen multilingual phenomena and domains, thereby assessing generalization beyond the released evaluation tasks.
@@ 4.5 Baselines
For the Multilingual track, we will be adapting our GPT-BERT and GPT-2 Small baselines to the new task setting. This will be done in a naive manner. Specifically, we will construct a 100M word dataset featuring English, Dutch and Chinese by subsampling an equal number of words from the respective dataset of each language, while taking into account each language’s Byte Premium (Arnett et al., 2024). We will then perform BPE tokenization on the resulting dataset. Training for GPT-BERT and GPT-2 Small will then proceed in the same fashion as in their Strict track counterparts.
@@ 4.6 Competition Submission
  • A datasheet describing the composition of the custom dataset and containing a download link (if not using a BabyLM-provided corpus)
@@ 5 FAQs
Moreover, we accept analysis and insightful findings on previous submissions or related topics and especially welcome evaluation that works well for small models but evaluates meaningful aspects. If you believe you know of an evaluation that we should use throughout the competition, please contact us.
@@
For evaluation purposes, we require that the model provides a function to score a sequence of words without the need for additional fine-tuning.
@@ 6 Organizing Committee
(Alphabetical by last name) Leshem Choshen, Ryan Cotterell, Mustafa Omer Gul, Jaap Jumelet, Tal Linzen, Aaron Mueller, Suchir Salhan, Raj Sanjay Shah, Alex Warstadt, and Ethan Wilcox. Feel free to contact members of the organizing committee at: leshem.choshen@mail.huji.ac.il, amueller@bu.edu, alexwarstadt@gmail.com

### v2

# BabyLM Turns 4 and Goes Multilingual: Call for Papers for the 2026 BabyLM Workshop https://babylm.github.io/

@@ Abstract
[…]
@@ 1 Introduction: BabyLM
The goals of BabyLM are to bring together multiple disciplines to answer an enduring question: how can a computational system learn language from limited inputs? Cognitive scientists investigate this question by trying to understand how humans learn their native language during childhood. Computer scientists tackle this question by attempting to build efficient machine-learning systems to accomplish this task. BabyLM aims to bridge these communities, encouraging the integration of insights from cognitive science into the design of more sample-efficient language models, while also using advances in language modeling architectures to generate new hypotheses and experimental paradigms for cognitive science.
@@ 3.1 Topics
  • Second language acquisition, bilingualism, or multilingualism.
@@
The workshop will have a theme this year: Going beyond English. Previous iterations of BabyLM have primarily focused on English, and with the introduction of the new Multilingual track we aim to inspire workshop submissions to investigate BabyLM-related topics with a focus on other languages. We hope the BabyBabelLM dataset can serve as a starting point for this, band we also encourage submissions introducing new resources that will promote progress in data-efficient language modeling across diverse linguistic settings. In particular, we welcome work that expands coverage to underrepresented and low-resource languages, explores cross-lingual transfer and multilingual training strategies, and examines how typological diversity, morphology, and script variation shape learnability under BabyLM constraints. We are also interested in benchmarks, evaluation protocols, and ethical data practices that support inclusive and culturally grounded language technologies.
@@ 4.2 Training Requirements
Similar to last year, we require the submission of results for models trained on a fixed amount of input (counting repeated exposures), in particular after at most 100M words for the Strict-small track and after at most 1B words for all other tracks. In most cases, this will mean after 10 epochs on the standard BabyLM corpora. However, because what counts as an epoch may differ across submissions, we instead quantify training by the number of whitespace-separated input words. For the Multilingual track, this value should be computed while taking each language’s Byte Premium into account (Arnett et al., 2024). While participants are welcome to train for longer and report this in their paper, we will only include the model checkpoints that follow these limitations in the leaderboard. Note that it is also allowed to submit a model that is trained on less data: the 100M and 1B word limits are an upper bound on data exposure. For the Strict and Strict-small tracks, the number of words seen by the submission model is considered to be the sum of the number of input words and generated tokens.
@@ 4.3 Updated Dataset
CHILDES (MacWhinney, 2000) Child-directed speech 15M 28.4M
British National Corpus (BNC), dialogue portion Dialogue 4M 7.6M
Project Gutenberg (children’s stories) (Gerlach and Font-Clos, 2018) Written English 13M 25.5M
OpenSubtitles (Lison and Tiedemann, 2016) Movie subtitles 10M 21.8M
Simple English Wikipedia Written Simple English 7M 15.3M
Switchboard Dialog Act Corpus (Stolcke et al., 2000) Dialogue <<1M 248K
@@
Recent work from last year’s Interaction Track found that BabyLMs trained on the provided Strict corpus can produce toxic generations and hate speech, which was a proven challenge for sustainable multi-turn generations between a BabyLM and a Teacher model (Salhan et al., 2025). Subsequent analysis by Trhlik et al. (2026) finds that the BabyLM corpus is more toxic and hateful than training corpora used in larger models such as BERT, with a comparatively higher presence of toxic and hateful sentences, even containing racist and sexualised training data, contrary to its child-aligned disposition.
@@
The dataset can be accessed on this Huggingface collection.
The datasets we provide serve as a starting point for training your models. We do, however, also allow participants to ‘swap’ (part of) the training data with other data sources of your choice, while adhering to the 10/100M word constraint. This allows for experimentation of more optimal training data; datasets that have been used in previous BabyLM editions have been made publicly available at the following link.
@@ 4.4 Evaluation
As in previous years, we will distribute an open-source evaluation pipeline, building on top of our repository from the 2025 challenge. Much of the evaluation will continue to be based on zero-shot probability comparisons of two text sequences. This year’s new track, Multilingual, will follow a similar logic to the existing evaluation for Strict. Specifically, participants will be evaluating their trained models on a combination of zero-shot and finetuning-based tasks for English, Dutch, and Chinese. More details about the evaluation pipeline and the set of tasks will be released subsequently. Similar to previous years, we will release a set of hidden tasks shortly before the submission deadline, to test the robustness of models on previously unseen multilingual phenomena and domains, thereby assessing generalization beyond the released evaluation tasks.
@@ 4.5 Baselines
For the Multilingual track, we will be adapting our GPT-BERT and GPT-2 Small baselines to the new task setting. This will be done in a naive manner. Specifically, we will construct a 100M word dataset featuring English, Dutch and Chinese by subsampling an equal number of words from the respective dataset of each language, while taking into account each language’s Byte Premium (Arnett et al., 2024). We will then perform BPE tokenization on the resulting dataset. Training for GPT-BERT and GPT-2 Small will then proceed in the same fashion as in their Strict track counterparts. We note again that this is intended to be a naive baseline. Participants can use different ratios of data between languages (or use custom datasets), experiment with tokenization and make contributions on architecture, training objectives, etc. as is the case with the Strict and Strict-small tracks.
@@ 4.6 Competition Submission
  • A datasheet describing the composition of the custom dataset and containing a download link (if not using a BabyLM-provided corpus). For the Multilingual track, please upload your training corpus containing your custom language mixture.
@@ 5 FAQs
Moreover, we accept analysis and insightful findings on previous submissions or related topics and especially welcome evaluations that work well for small models but evaluate meaningful aspects. If you believe you know of an evaluation that we should use throughout the competition, please contact us.
@@
For evaluation purposes, we require that the model provide a function to score a sequence of words without the need for additional fine-tuning.
@@ 6 Organizing Committee
(Alphabetical by last name) Leshem Choshen, Ryan Cotterell, Mustafa Omer Gul, Jaap Jumelet, Tal Linzen, Aaron Mueller, Suchir Salhan, Raj Sanjay Shah, Alex Warstadt, and Ethan Gotlieb Wilcox. Feel free to contact members of the organizing committee at: leshem.choshen@mail.huji.ac.il, amueller@bu.edu, alexwarstadt@gmail.com